# ACEBench: Who Wins the Match Point in Tool Learning?

### v2

@@ 3.2 Eval
In this section, we introduce the evaluation process, where different evaluation metrics and methods are applied based on the various data types.
@@ 3.2.1 Normal Evaluation
As shown in the left of Figure 5, we evaluate Normal Data by comparing the model’s function call outputs with the ground truth. The inference prompt used for Normal Data is shown in Appendix B.1, which guides the model’s reasoning process in generating accurate responses. We conduct the evaluation using AST parsing, beginning with the matching of the function name to the target function name to ensure exact correspondence. Next we match the parameter types to verify their consistency with the target. Finally we compare the parameter values to confirm their accuracy. For cases with multiple possible correct answers, we use a candidate answer pool, and a match with any one of the candidates is considered correct. This step-by-step evaluation allows us to progressively assess the consistency between the model’s output and the ground truth, providing a clear and accurate measure of the model’s performance in tool usage.
Metrics The Normal Data is primarily evaluated using the Accuracy, where a score of 1 is assigned for a complete match, and a score of 0 is given for a mismatch.
@@ 3.2.2 Special Evaluation
For the Special Data, the inference prompt in Appendix B.2 guides the model in identifying these issues and outputting the results in the specified format. And as shown in the middle of Figure 5, the evaluation of Special Data can be categorized into three distinct scenarios.
The first scenario involves the absence of necessary parameters. It occurs when the instruction fails to include all the parameters required for execution. Only when the model correctly identifies all required parameters, the Accuracy is assigned a value of 1.
The second scenario pertains to erroneous parameter representations. Many functions enforce strict conventions for parameter names or values, or they define a range of acceptable values. If any parameter name or value in the instruction deviates from these conventions, the instruction is deemed erroneous. If the model successfully identifies and points out the incorrect parameter, the Accuracy is 1. If the model fails to recognize the erroneous parameter, the Accuracy is 0.
The third scenario occurs when the instruction exceeds the function’s processing capabilities. This situation arises when the task or request specified in the instruction cannot be processed by the current function, either due to the function’s inherent limitations or the complexity of the parameter combinations involved. If the model correctly identifies the issue and indicates that the function cannot process the request, it is considered a success, and the Accuracy is 1. If the model fails to recognize the issue, the Accuracy is 0.
@@ 3.2.3 Agent Evaluation
As shown in the right of Figure 5, we evaluate the Agent Data by assessing the model’s ability to use tools in real-world interactions with humans, focusing on both result and process evaluation. Based on various real-world scenarios, we abstract them into distinct instances, each of which defines functions that meet the specific needs of the scenario.
There are two evaluation metrics: End Accuracy and Process Accuracy.
  • End Accuracy (Accuracy) is evaluated by comparing the instance attributes of the corresponding class with the target. If all attributes match exactly, the accuracy is 1; if they do not match completely, the accuracy is 0. This metric reflects the overall correctness of the result.
  • Process Accuracy is based on the consistency between the actual function call process and the ideal function call process. It is expressed as $\frac{n}{m}$, where $m$ represents the ideal function call process, and $n$ represents the degree of match between the actual function call process and the ideal process. This metric evaluates how closely the actual execution matches the expected sequence of steps.
@@ 3.2.4 Overall Accuracy
The Overall Accuracy is computed as a weighted sum of the accuracies for the Normal, Special, and Agent data types, where the weights are determined by the square roots of their respective sample sizes. The details can be found in Appendix D.
Figure 6: Example of Agent Data. The left shows the dialogue process, highlighting the interaction between the model and the user. The right presents the evaluation process, including result evaluation and process evaluation.
@@ 4 Experiments
Experimental Setup:
In our evaluation, we examine eight closed-source LLMs, including the GPT-4 series Achiam et al. (2023), DeepSeek-Chat Liu et al. (2024a), Qwen-Max Yang et al. (2024), Gemini-1.5-Pro Team et al. (2024), Claude-3.5-Sonnet Anthropic (2024), and Doubao-Pro-32K ByteDance (2025). Additionally, a wide range of open-source language models are assessed, such as the Qwen2.5 series Yang et al. (2024), Llama3 series Dubey et al. (2024), Phi-3-Mini Abdin et al. (2024), and DeepSeek-Coder-V2 Zhu et al. (2024). Furthermore, four tool-learning-enhanced models were evaluated: Hammer2.1-3B, Hammer2.1-7B Lin et al. (2024), xLAM-7B-R Liu et al. (2024c), and Watt-Tool-8B Watt-AI (2024).
@@ 4.1 Main results and analysis
The comprehensive experimental results for both the Chinese and English datasets are presented in Table 2, with detailed results for each language provided in Appendix F. These findings highlight an important observation: the overall best performance still comes from closed-source models like the GPT-4 series, but the performance gap between certain open-source models, such as Qwen2.5-Coder-32B-Instruct and Qwen2.5-72B-Instruct, and existing closed-source models is narrowing. In addition, we can draw the following important conclusions:
Normal Data Performance and Model Size Limitations: As shown in Table 2, most models perform well on Normal data, indicating their effective acquisition of fundamental tool invocation abilities during training. This strong performance stems from the models’ focus on mastering basic function calling capabilities. However some open-source models show relatively weaker performance, which can be attributed to the limitations imposed by their smaller model sizes, restricting their capacity to handle more complex tasks and interactions effectively.
Decline in Fine-Tuned Models’ Performance on Special Data: As shown in Table 3, models fine-tuned on specific datasets, such as Hammer2.1-7B Lin et al. (2024), Watt-Tool-8B Watt-AI (2024), and xLAM-7b-r Liu et al. (2024c), exhibit a significant decline in performance on the Special dataset. This decline can primarily be attributed to the fact that while fine-tuning improves the model’s performance on specialized tasks, it leads to a loss of general instruction-following ability. As a result, the model’s accuracy in handling general instructions decreases after fine-tuning on special data.
Table 3: The Accuracy of Different Models on Special Data
Model Lose Error Exceed Summary
Closed-Source Large Language Models
GPT-4o 0.86 0.94 0.99 0.93
Gemini-1.5-Pro 0.67 0.76 0.93 0.79
Claude-3-5-Sonnet 0.64 0.72 0.96 0.78
Qwen-Max 0.62 0.68 0.92 0.74
Open-Source Large Language Models
Qwen2.5-72B-Instruct 0.57 0.75 0.95 0.76
Qwen2.5-7B-Instruct 0.26 0.36 0.79 0.47
Hammer2.1-7b 0.03 0.18 0.23 0.15
Watt-Tool-8B 0.07 0.01 0.10 0.06
xLAM-7b-r 0.01 0.03 0.04 0.03
Performance Limitations of Large Models in Agent Data: As shown in Table 4, most models exhibit an end accuracy of less than 50% on Agent data tasks. This phenomenon can be attributed to the fact that completing a task in a dynamic environment, simulating real-world multi-turn interactions, requires more than just the model’s ability to perform individual tool operations. It also demands that the model effectively integrate contextual information during tool usage and consider the interdependencies between tool calls, making the task more challenging.
Table 4: Results of Different Models on Agent Data
@@
EndAcc ProcessAcc EndAcc ProcessAcc
GPT-4-Turbo 0.50 0.66 0.85 0.96
Deepseek-Chat 0.32 0.55 0.38 0.53
Claude-3-5-Sonnet 0.22 0.42 0.58 0.77
DouBao-Pro-32k 0.20 0.46 0.30 0.48
Qwen2.5-7B-Instruct 0.13 0.28 0.13 0.16
Hammer2.1-7b 0.09 0.34 0.25 0.43
Radar Analysis of Top-3 Open-source and Closed-source Models: Through the radar chart analysis of the top three open-source and top three closed-source models, as shown in the Figure 7, we draw the following conclusions: (1) Performance in basic ability tasks is similar: These top models perform closely in normal tasks, all achieving high levels. Among them, the GPT-4 series models lead in most tasks, demonstrating strong intelligence. Overall, closed-source models generally outperform open-source models, thanks to their more refined internal optimizations and possibly larger scales. (2) Significant model differences in mid-to-high-level tasks: There is a noticeable performance gap between the models in tasks such as Special and Agent. These tasks require higher-level abilities: Special tests the model’s capability to handle abnormal situations, while Agent require strong multi-turn interaction and reasoning capabilities.
Figure 7: Performance of Top-3 Closed/Open-source models.
@@ 4.2 Error Analysis
Error Analysis of Normal Data: By analyzing the error type distribution across different model series on the Normal data, as shown in Figure 8, we observe that Wrong Value Errors dominate across all models. This highlights the models’ difficulty in accurately generating specific values, which likely stems from insufficient contextual understanding and the complexity of numerical distributions. AST Parsing Errors follow as the second most common error type, indicating significant room for improvement in generating code that adheres to predefined formats and syntactic rules. These challenges may result from inconsistencies in training data and the models’ limited ability to learn rule-based generation effectively. In contrast, Wrong Function and Wrong Type Errors occur far less frequently, showing that the models excel in matching function calls and handling data types. Overall, while the models demonstrate strong function invocation capabilities, further advancements are needed in numerical generation and compliance with formatting rules. Specific error examples for Normal data can be found in Appendix C.1.
@@
Error Analysis of Special Data: Through the analysis of special type data, we identified two main types of model errors( specific examples can be found in Appendix C.2): The first type is "Error Blindness", which refers to the model’s complete failure to detect issues in the user’s instructions or its inability to identify problems according to the prompt’s formatting requirements. This highlights significant flaws in the model’s error detection capabilities or its general instruction-following ability. The second type is "Vague Error Reporting," where the model detects the problem but provides unclear feedback. For example, the model might indicate that there is an issue, but fails to specify which parameter values are incorrect or what critical information is missing. As shown in Table 5,our experimental results indicate that the majority of errors in special type scenarios are caused by "Error Blindness", highlighting a critical gap in the model’s problem detection capabilities. This suggests that the model needs to learn not only simple tool invocation but also how to identify corresponding issues under imperfect instructions.
Table 5: Error Blindness and Vague Error Reporting Across Models
Model Error Blindness Vague Error Reporting
@@
Meta-Llama-3.1-8B-Instruct 145 6
Qwen2.5-Coder-3B-Instruct 168 9
@@
Error Analysis of Agent Data: We evaluate the Agent type data to assess the LLM ability to effectively perform function calls in interactive scenarios. In this process, the model must not only demonstrate accurate learning of function calls but also strictly adhere to predefined scene rules. In Appendix B.3, we present the system prompt for one of these scenarios. Our analysis identifies three primary causes of Agent errors. First, function call errors occur when the model fails to select the appropriate function or provide parameters that do not meet the required specifications, reflecting a lack of understanding of function calling capabilities and parameter constraints. Second, rule violations arise when the model disregards predefined scene rules, skipping necessary steps or breaking key task logic, which highlights deficiencies in its comprehension and execution of these rules. Finally, information mismanagement results from the model’s inability to correctly record or process contextual information during multi-turn interactions, leading to outputs that diverge from expectations. As shown in Figure 6, we illustrate an error caused by missing information.
@@ 4.3 Further Analysis
Scaling Law: We evaluated the performance of Qwen2.5-Coder (7B, 14B, 32B) and Qwen2.5-Instruct (3B, 7B, 14B, 32B, 72B) on the ACEBench dataset. The experimental results shown in the Figure 9 demonstrate that as the model size increases, performance improves significantly across various tasks, with particularly strong results on high-complexity tasks. This suggests that increasing the parameter size enables the model to achieve stronger task-solving and generalization capabilities.
@@
Impact of Code Training: Based on the experimental results in Figure 9, we find that the score of Qwen2.5-Coder-7B is lower than the non-coder version, while the scores of Qwen2.5-Coder-14B and Qwen2.5-Coder-32B are higher than those of the non-coder versions of the same size. This indicates that as the model size increases, code tuning becomes more impactful in improving model performance. For larger models (such as 14B and 32B), code tuning significantly enhances tool usage, especially in tasks that follow predefined rules and steps. However, for smaller models (such as 7B), despite code tuning, the differences between models are minimal, suggesting that smaller models have limited capacity and cannot fully leverage the advantages of code tuning.
@@ 5 Conclusion
This paper introduces ACEBench, a comprehensive tool learning benchmark designed to evaluate the function calling capabilities of large language models. It addresses key limitations of existing evaluation systems, such as the lack of multi-turn dialogue assessments in real-world scenarios, the absence of fine-grained evaluations for parameter-type function calls, and the high costs associated with using large models for evaluation. ACEBench provides multi-perspective and multi-scenario evaluations, offering a comprehensive assessment of large models’ function calling capabilities in different contexts. The experimental results show that: (1) Some models fine-tuned on specific function calling datasets are capable of handling general function calls, but perform poorly when faced with complex or imperfect instructions, indicating a lack of generalization ability; (2) Code capability improves the function calling performance of large models, especially in tasks that follow predefined rules. Through extensive experiments, we demonstrate the effectiveness of ACEBench in providing deeper insights into the function calling abilities of various models.
Appendix A Examples of Dataset
A.1 Normal Examples

### v3

@@ 3.2 Eval
In this section, we introduce the evaluation process, where different evaluation methods are applied based on the various data types. The inference prompt for evaluation can be found in Appendix D.
@@ 3.2.1 Normal Evaluation
As shown in the left part of Figure 5, we evaluate Normal Data by comparing the model’s function call outputs with the ground truth. We evaluate using AST parsing, starting by matching the function name to the target for exact correspondence, followed by verifying the consistency of parameter types, and finally confirming the accuracy of the parameter values. For cases with multiple possible correct answers, we use a candidate answer pool, and a match with any one of the candidates is considered correct. The Normal Data is evaluated using Accuracy, where a score of 1 indicates a complete match and 0 indicates a mismatch.
@@ 3.2.2 Special Evaluation
As shown in the middle part of Figure 5, the core goal of evaluating Special Data is to determine whether the model can accurately identify and point out the issue. Specifically, when the instruction is missing required parameters, the model needs to identify and indicate the missing parameter; when there is an error in the parameter, the model needs to correctly identify the error parameter; and when the task is irrelevant to the functions, the model needs to recognize that the function cannot process the request. In any of these cases, if the model correctly identifies and addresses the issue, the Accuracy is 1; otherwise, the Accuracy is 0.
@@ 3.2.3 Agent Evaluation
As shown in the right part of Figure 5, we evaluate the Agent Data by assessing the model’s ability to use tools in real-world interactions with humans, focusing on both result and process evaluation. There are two evaluation metrics:
End-to-End Accuracy is evaluated by comparing the instance attributes of the corresponding class with the target. If all attributes match exactly, the Accuracy is 1; otherwise, the Accuracy is 0.
Process Accuracy is determined by the consistency between the actual function call process and the ideal process. It is expressed as $\frac{n}{m}$, where $m$ represents the ideal function call process, and $n$ represents the degree of match between the actual and ideal processes.
@@ 3.2.4 Overall Accuracy
The Overall Accuracy is computed as a weighted sum of the accuracies for the Normal, Special, and Agent data types, where the weights are determined by the square roots of their respective sample sizes. The details can be found in Appendix F.
@@ 4 Experiments
Experimental Setup. In our evaluation, we examine seven closed-source LLMs, including the GPT-4 series Achiam et al. (2023), Qwen-Max Yang et al. (2024), Gemini-1.5-Pro Team et al. (2024), Claude-3.5-Sonnet Anthropic (2024), and Doubao-Pro-32K ByteDance (2025). Additionally, a wide range of open-source language models are assessed, such as the Qwen2.5 series Yang et al. (2024), Llama3 series Dubey et al. (2024), Phi-3-Mini Abdin et al. (2024), Deepseek-V3Liu et al. (2024a), and DeepSeek-Coder-V2 Zhu et al. (2024). Furthermore, four tool-learning-enhanced models were evaluated: Hammer2.1-3B, Hammer2.1-7B Lin et al. (2024), xLAM-7B-r Liu et al. (2024c), and Watt-Tool-8B Watt-AI (2024).
@@ 4.1 Main results and analysis
The comprehensive experimental results for the Chinese and English datasets are presented in Table 2, with detailed results for each language provided in Appendix F.1. We can draw the following important conclusions:
Performance Variations Between Top Open-source and Closed-source Models. The overall best performance still comes from closed-source models like the GPT-4 series, but the performance gap between certain open-source models, such as Qwen2.5-Coder-32B-Instruct and Qwen2.5-72B-Instruct, and existing closed-source models is narrowing. And as shown in Figure 6, the performance of these top models is generally similar in basic ability tasks, with the GPT-4 series models leading in most tasks, demonstrating strong intelligence. However, closed-source models outperform open-source models due to more refined internal optimizations and potentially larger scales. A noticeable performance gap emerges in mid-to-high-level tasks, such as Special and Agent, as these tasks require higher-level abilities, such as handling abnormal situations and strong multi-turn interaction and reasoning capabilities.
Figure 6: Performance of Top-3 Closed/Open-source models.
Loss of Generalization in Fine-Tuned Models. As shown in Table 3, models fine-tuned on specific datasets, such as Watt-Tool-8B Watt-AI (2024), xLAM-7B-r Liu et al. (2024c) and Hammer2.1-7B Lin et al. (2024), exhibit a significant decline in performance on the Special dataset. This decline can primarily be attributed to the fact that while fine-tuning improves the model’s performance on specialized tasks, it leads to a loss of general instruction-following ability.
Table 3: The Accuracy of Different Models on Special Data (%).
Model Incomplete Error Irrelevant
Llama-3.1-8B-Instruct 29.0 20.0 14.0
Watt-Tool-8B 7.0 1.0 10.0
Qwen2.5-7B-Instruct 26.0 36.0 79.0
xLAM-7B-r 1.0 3.0 4.0
Llama-3.2-3B-Instruct 13.0 12.0 3.0
Hammer2.1-3B 0.0 3.0 0.0
Performance Limitations of Large Models in Complex Tasks. As shown in Table 4, most models exhibit an end accuracy of less than 50% on Agent data tasks. This phenomenon can be attributed to the fact that completing a task in a dynamic environment, simulating real-world multi-turn interactions, requires more than just the model’s ability to perform individual tool operations. It also demands that the model effectively integrate contextual information during tool usage and consider the interdependencies between tool calls, making the task more challenging.
Table 4: Performance Evaluation of Different Models on Agent Data: PA represents Process Accuracy, EA represents End-to-End Accuracy (%).
@@
EA PA EA PA
GPT-4-Turbo 50.0 66.06 85.0 89.5
DeepSeek-V3 31.5 54.5 37.5 53.0
Claude-3-5-Sonnet 21.5 41.5 57.5 76.5
DouBao-Pro-32k 20.0 45.5 30.0 47.5
Qwen2.5-7B-Instruct 15.0 28.0 12.5 15.5
Hammer2.1-7B 8.5 33.5 25.0 42.5
@@ 4.2 Error Analysis
Error Analysis of Normal Data. As shown in Figure 7, we observe from the error type distribution on Normal data that param value error dominate across all models. This highlights the models’ difficulty in generating specific values, likely due to limited contextual understanding and the complexity of numerical distributions. Output format error is the second most common, suggesting room for improvement in generating code that follows predefined formats and syntactic rules. These issues may stem from inconsistencies in training data and the models’ limited ability to learn rule-based generation. In contrast, function name and param type errors are less frequent, indicating that the models excel in matching function calls and handling data types. While the models show strong function invocation abilities, further improvements are needed in numerical generation and format compliance. Specific error examples for Normal data can be found in Appendix E.1.
@@
Error Analysis of Special Data. As shown in Table 5, we identified two main types of model errors: The first type is "Error Detection", which refers to the model’s complete failure to detect issues in the user’s instructions or its inability to identify problems according to the prompt’s formatting requirements. The second type is "Error Correction," where the model detects the problem but provides unclear feedback. For example, the model might indicate that there is an issue, but fails to specify which parameter values are incorrect or what critical information is missing. Results show that most errors in special-type scenarios are caused by "Error Detection", highlighting a critical gap in the model’s problem-detection capabilities. This suggests that the model needs to learn not only simple tool invocation but also how to identify corresponding issues under imperfect instructions. Specific error examples for Special data can be found in Appendix E.2.
Table 5: Error type distribution across different model series on Special Data.
Model Error Detection Error Correction
@@
Llama-3.1-8B-Instruct 145 6
@@
Error Analysis of Agent Data. Our analysis identifies three primary causes of Agent errors. First, function call errors occur when the model fails to select the appropriate function or provide parameters that do not meet the required specifications, reflecting a lack of understanding of tool-use capabilities and parameter constraints. Second, rule violations arise when the model disregards predefined scene rules, skipping necessary steps or breaking key task logic, highlighting deficiencies in its comprehension and execution. Finally, information mismanagement results from the model’s inability to correctly record or process contextual information during multi-turn interactions, leading to outputs that diverge from expectations. As shown in Figure 20, we illustrate an error caused by missing information.
@@ 4.3 Further Analysis
Scaling Law. We evaluated the performance of Qwen2.5-Coder (7B, 14B, 32B) and Qwen2.5-Instruct (3B, 7B, 14B, 32B, 72B) on the ACEBench dataset. The experimental results shown in Figure 8 demonstrate that performance improves significantly across various tasks as the model size increases, with particularly strong results on high-complexity tasks. This suggests that increasing the parameter size enables the model to achieve stronger task-solving and generalization capabilities.
Impact of Code Training. Based on the experimental results in Figure 8, we find that the score of Qwen2.5-Coder-7B is lower than the non-coder version, while the scores of Qwen2.5-Coder-14B and Qwen2.5-Coder-32B are higher than those of the non-coder versions of the same size. This indicates that code tuning becomes more impactful in improving model performance as the model size increases. For larger models (such as 14B and 32B), code tuning significantly enhances tool usage, especially in tasks that follow predefined rules and steps. However, for smaller models (such as 7B), despite code tuning, the differences between models are minimal, suggesting that smaller models have limited capacity and cannot fully leverage the advantages of code tuning.
@@ 5 Conclusion
This paper introduces ACEBench, a comprehensive tool-use benchmark designed to evaluate the tool-use capabilities of Large Language Models (LLMs). It addresses key limitations of existing evaluation benchmarks, such as the lack of multi-turn dialogue assessments in real-world scenarios, the absence of fine-grained evaluations for parameter-type function calls, and the high costs associated with using large models for evaluation. The experimental results indicate that models fine-tuned on specific tool-use datasets to struggle with generalization when faced with complex or imperfect instructions, and code capabilities enhance the tool-use performance of large models. Through extensive experiments, we demonstrate the effectiveness of ACEBench in providing deeper insights into the tool-use abilities of various models.
Limitations
We acknowledge several limitations in our evaluation of ACEBench for assessing the tool-use capabilities of large language models. Firstly, while our test data is generated by large language models and various measures have been taken to ensure its authenticity and diversity, a gap remains when compared to data from real-world applications. This discrepancy may impact the evaluation of the model’s performance in real-world scenarios. Secondly, for the Agent data, the design of evaluation scenarios relies on manual construction, which somewhat limits the diversity and coverage of the evaluation framework.
Appendix A Detailed Descriptions of Test Cases
A.1 Data Categories Description
Normal Data
The Normal Data consists of fixed question-answer pairs, where each question corresponds to a correct function call. It is categorized into the following categories: Single Turn, Multi Turn, Similar API, Preference and Atom. Data examples can be found in Appendix C.
Multi Turn: There are multiple interactions between the user and the assistant. The conversation can be categorized into two types: (switch) The conversation progresses by changing topics. (adjust) The conversation evolves by refining or modifying the original question.
Similar APIs: The candidate APIs exhibit significant similarity, particularly in their focus on the same topic. This similarity presents a challenge for the assistant, requiring it to effectively distinguish between the APIs and accurately select the most appropriate ones.
Special Data
The Special Data refers to situations where the model is unable to resolve the problem posed in the instruction using the candidate functionsWang et al. (2024c). It is categorized into the following categories: Lose param, Error param, and Exceed Function Ability. Data examples can be found in Appendix C.2.
Incomplete: Refers to situations where the key information required for the function call is missing in the query, such as the absence of "required" parameters.
Error: Refers to situations where the instruction contains parameters or names that do not meet the required format or constraints, such as matching a specific pattern or being selected from a predefined list, causing the function call to fail.
Irrelevant: Refers to situations where the instruction exceeds the function’s capabilities, meaning none of the candidate functions can resolve the issue.
Agent Data
Agent Data refers to situations where completing a task in an environment built based on real-world scenarios typically requires multiple steps. In such cases, an advanced language model simulates the user role, replicating interactions found in real-world scenarios. This approach aims to evaluate the model’s ability to perform in complex interactions. We define a "multi-step" scenario as one in which the user appears only once in the entire conversation, while a "multi-turn" scenario is one where the user appears multiple times throughout the conversation. Data examples can be found in Figure C.3.
Appendix B The Construction of Data
B.1 Agent Data
Second, implement logical abstraction for each function through code, designing core processes such as user authentication, cost calculation, payment processing, and information recording. This ensures the code logic’s scalability and robustness while comprehensively addressing exception handling.
B.2 Special Data
Incomplete
The special data which has incomplete instructions mainly refers to situations where key information is missing from the user’s instructions, causing the function to be called incorrectly. Our main prompt for constructing incomplete data is shown in  10. And we can extract the corresponding data and reference answers (missing data) from the generated dialogue.
{mdframed}
Please refer to the example and continue the dialogue based on the given tool definition and the beginning of the conversation. The requirements are as follows:
1 The user’s request is meaningful, requiring the use of one tool, and the tool will be called once.
2 Calling the tool to fulfill the user’s request still lacks one or more required parameters .
3 The assistant’s first response should be formatted as: [Thought] Missing parameters: Key1|Key2|…[Response] (follow-up question)". Replace Key1, Key2, etc., with specific parameter names, and replace the follow-up question with natural language content.
Here is an example for reference
<tool_definition>
"name": "calculate_triangle_area",
"description": "Given the base and height of a triangle, calculate its area.",
"required": ["base", "height"]
<dialogue_example>
[User]: I want to calculate the area of a triangle with a base of 5 and a height of 10.
[Assistant]: [Thought]  Missing parameters: base|height
[Response]:  Please provide both the base and the height of the triangle.
[User]: The base is 5, and the height is 10.
[Assistant]: <tool_usage>calculate_triangle_area|{"base": 5, "height": 10}</tool_usage>
Here is the dialogue continuation you need to write:
<Tool Definition>
{tool_definition}
Figure 10: Example of Incomplete Data construction.
Next, we can convert the conversation we obtained into data. Here is a specific example.
{mdframed}
candidate function
  "name": "book_flight",
  "description": "Flight booking tool...",
  "arguments": {},
  "required": ["from_city_name", "to_city_name", "depart_date"]
obtained dialouge
<user> I plan to travel to Beijing. Help me book a flight to Beijing on June 1st.
</user> [Thought] Missing parameters: from_city_name
[Response] May I know your departure city you want to book?
<user> Oh, Im in Shenzhen, economy class X
converted data
[question] I plan to travel to Beijing. Help me book a flight to Beijing on June 1st.
[answer] Miss necessary parameter (from_city_name) from the api (book_flight)
Error
The special data which has error instructions mainly refers to situations where the instruction contains parameters or names that do not meet the required format or constraints
{mdframed}
Please refer to the example and continue the dialogue based on the given tool definition and the beginning of the conversation. The requirements are as follows:
1 The user’s request is specific and meaningful, requiring the use of one tool, and the tool should be called once.
2 The user’s first question contains incorrect parameters (i.e., it does not conform to the function definition’s pattern or format requirements).
3 The assistant’s first reply should accurately identify the error and ask the user for more precise parameters, formatted as: ’Thought Error in parameter: Key1 Response (Follow-up question content)’. Replace ’Key’ with the specific parameter name and provide a natural, real-language follow-up question."
Here is an example for reference
<tool_definition>
{"name": "TennisTeamDetails",
  "parameters": {
    "properties": {
      "team_name": {......
[…]
<dialogue_example>
<user> Can you retrieve the details of the tennis team named ’Team@1234’?
</user> [Thought] Error Parameter: ’Team@1234’
[Response]  The team name you provided does not meet the naming criteria.
<user> Oh, I made a mistake. The team name is TeamABC.
</user> <tool_usage>TennisTeamDetails|{"team_name": "TeamABC"}</tool_usage>"
Here is the dialogue continuation you need to write:
<Tool Definition>
{tool_definition}
Figure 11: Example of Incomplete Data construction.
Next, we can convert the conversation we obtained into data. Here is a specific example.
{mdframed}
candidate function
{ "name": "FootballTeamDetails",
  "description": "Retrieve detailed information about a football team by its name.",
  "parameters": {
      "team_name": {......
[…]
obtained dialouge
<user> I want to know information about the football team football\$156.
</user> [Thought] Error Parameter: football\$156.
[Response] The team name you provided does not meet the naming criteria.
<user> ......
converted data
[question] I want to know information about the football team footbaoo\$156.
[answer] There is incorrect value (football\$156) for the parameters (team_name).
Irrelevant
The construction method for Irrelevant data is similar to that of Normal data, where we simply remove the correct API from the candidate APIs of the data.
Appendix C Examples of Dataset
C.1 Normal Examples